# Knolling Bot: Learning Robotic Object Arrangement from Tidy Demonstrations

### v2

@@ 5. Entropy Loss ($L_{\text{entropy}}$)
This loss function encourages diversity in the predictions by measuring the entropy across the probabilities of the five Gaussian distributions for each prediction. The entropy loss is defined as:
@@ 6. Weights of Loss Components
We assigned each loss component a weight to balance the contributions of each loss during the training process. We used MSE Min loss and overlapping loss to evaluate the knolling performance.
III-D Knolling Pipeline
Our knolling pipeline comprises three modules: a knolling model, a visual perception model, and the robot arm controller for executing the pick-and-place task.
The visual perception model is a customized YOLO v8 model [16], fine-tuned to predict the information of the objects, including the segmentation, color, and category. We post-process the segmentation to a rectangular box and then extract the object state, including the center position, orientation, and dimensions. Detailed evaluation results and confusion matrices for the visual perception model in both simulated and real-world environments are provided in the supplementary materials. The data required for training our models is generated within the Pybullet simulation environment for pre-training and collected in the real world for fine-tuning. To prepare our Visual Perception Model for real-world deployment and address the simulation-to-reality gap, we apply the visual domain randomization technique to our data collection process [49]. This technique helps in enhancing the model’s adaptability and robustness by introducing variations and uncertainties that mimic real-world conditions. We manipulate several factors within the simulation environment, such as brightness and ground textures. Additionally, we also vary the appearance of the objects with different colors. Using this approach, we train our model to disregard extraneous noise and focus on essential features.
IV Experiments
IV-A Qualitative Evaluations in the Simulation
IV-A1 Handling Varying Object Quantities
In real-world scenarios, the number of objects on a surface can vary significantly. To assess our model’s ability to handle this variability, we conducted experiments with diverse object counts ranging from 2 to 10 items. Our model, leveraging its autoregressive nature, successfully generated tidy arrangements across all object quantities, as shown in Fig. 4. This adaptability to varying input sizes is a crucial aspect of our approach, enabling it to generalize to different scenarios without explicit retraining.
IV-A2 Generating Diverse Solutions Based on Preferences
Tidiness preferences can differ significantly among individuals, leading to multiple valid solutions for the same set of objects. To capture this diversity, our model decouples preferences from the input data (which consists solely of object dimensions) and instead encodes preferences through the order of the input sequence. By altering the input order based on criteria such as object category, color, or size, our model can generate distinct tidy arrangements tailored to specific preferences, as illustrated in Fig. 5. This flexibility to incorporate user preferences without modifying the model architecture is a key strength of our approach.
Figure 5: Knolling with preferences. This figure shows distinct layouts achieved by controlling the input order based on different criteria: 1. Categories: Place the same functionalities object together, 2. Shape: Big area priority, 3. Color: Group the objects that have the same colors.
IV-B Real-World Knolling Experiments
To demonstrate the practical applicability of our framework, we deployed our trained knolling model on a 5-DoF robotic arm (WidowX 200) equipped with an Intel Realsense D435 camera for visual perception. We randomly placed 6-10 boxes of varying sizes and colors within the robot’s workspace, simulating cluttered tabletop scenarios.
The integrated pipeline (Fig. 2) first captures an overhead image of the cluttered scene, which is processed by our visual perception module to detect objects, perform segmentation, and extract their dimensions and poses. The knolling model then predicts the target positions for each object based on their dimensions. Finally, the robotic arm controller executes smooth pick-and-place operations, guided by the predicted target positions and current object poses, to realize the tidy arrangements.
Fig. 6A showcases real-world knolling tasks performed on different box configurations, demonstrating our system’s capability to organize diverse object sets. Fig. 6B further illustrates the system’s adaptability to varying object quantities, successfully knolling scenes with 6, 8, and 10 objects. Additionally, Fig. 6C highlights the generation of distinct tidy arrangements for the same set of objects, reflecting different user preferences based on category, color, or shape. We encourage readers to view our supplementary video for a comprehensive visualization of these real-world knolling experiments.
IV-C Quantitative Evaluation and Ablation Study
@@
TABLE I: Comparisons between our methods and baselines
@@
TABLE II: Ablation study
@@
Two baseline models were established to enable a relative performance comparison to Our Method (OM). The first baseline model utilized a Multilayer Perceptron (MLP) architecture, and the second baseline was designed based on the Long Short-Term Memory (LSTM) model. We use LSTM as a baseline because knolling can also be viewed as a sequence prediction problem where the arrangement of objects depends on their sequence. LSTMs have been extensively used for various sequence prediction tasks and have shown good performance. Thus, LSTMs serve as a more challenging baseline to benchmark the performance of our transformer-based model. Moreover, comparing our model with LSTM helps to highlight the benefits of using self-attention and auto-regression in the transformer for handling variable input and output sizes and multi-label problems. Just as we did with our model, we measured the L1 distance between the actual and predicted positions for each baseline model, with the results detailed in Table I. As a commitment to a fair evaluation, we have ensured that each model utilizes a similar amount of parameters: our model incorporates 87,458 parameters, the LSTM baseline uses 86,858 parameters, and the MLP baseline operates with 87,788 parameters. Our knolling model consistently outperforms the MLP and LSTM baselines in terms of the L1 error. This superiority of our model is evident in all parameters: mean L1 error, standard deviation, and minimum and maximum error. These results validate the effectiveness of our transformer-based model, which employs self-attention and auto-regression in handling varying input sizes and multi-label problems in knolling tasks.

### v3

@@ 3. MSE Min Loss ($L_{\text{MSE Min}}$)
In a multi-target task, relying solely on MSE could lead the model to predict average positions in the presence of ambiguity, which is not desirable. The MSE Min Loss addresses this by allowing the model to choose the prediction (from among multiple Gaussian components) that is closest to the actual target position.
@@ 5. Entropy Loss ($L_{\text{entropy}}$)
This loss function encourages diversity in the predictions by measuring the entropy across the probabilities of the J Gaussian components for each prediction. The entropy loss is defined as:
@@ 6. Weights of Loss Components
Each loss component was assigned a specific weight to balance their contributions during training, with the aim of optimizing knolling performance. The weights were determined empirically, through iterative experimentation and performance evaluation, to ensure the model adequately minimized both positional errors and object overlap.
Appendix B Iterative Optimization for Knolling Dataset Generation
While the rule-based arrangement method presented here provides a systematic approach to generating a knolling dataset, it also has inherent limitations. Rule-based methods excel at articulating specific, human-defined rules of neatness—such as alignment, spacing, and categorization by color or type—that are relatively straightforward to parameterize. However, much of neatness is inherently subjective and nuanced. Although humans can often recognize whether an arrangement is “neat” or “not neat,” they may struggle to articulate the underlying principles that define these notions. This subjectivity poses a challenge for rule-based systems, which can only operate within explicitly defined rules and fail to capture the subtleties of more subjective or abstract concepts of neatness.
Defining an exhaustive set of rules to account for all possible tidiness concepts is infeasible, as the number of potential layouts grows exponentially with object variety, user preferences, and spatial constraints. As a result, rule-based or optimization methods alone cannot cover all valid solutions. Instead, we take the opposite approach, learning a generalizable representation of tidiness from demonstrations through self-supervised learning.
While rule-based optimization is valuable for data generation, it is often computationally expensive and inefficient for real-time object arrangement. In contrast, the learned knolling model operates in a single forward pass, making predictions significantly faster than iterative optimization. This efficiency enhances scalability, enabling practical deployment in real-world robotic applications.
The rule-based optimization bootstraps the learning process, providing structured training examples to establish an initial understanding of tidiness. However, it is not a substitute for data-driven method. The model leverages this structured data as a foundation but learns beyond it, capturing nuanced spatial relationships and placement strategies that a purely optimization-based approach could never explicitly define.
B.1 Objects Used in the Knolling Task
we provide the following details regarding our dataset. Rectangular Objects: A subset of the dataset consists of randomly generated rectangular objects, with their length and width sampled from a predefined range. These synthetic objects serve as fundamental elements in training the model to generalize tidiness principles. Some of these generated objects were physically printed and used in real-world experiments to validate the learned representations. Laboratory Items: In addition to the rectangular objects, we incorporated commonly used laboratory items such as batteries, erasers, and electronic components. These objects were selected to introduce variation in shape and appearance. To facilitate color preference learning, we 3D-printed different versions of these objects in multiple colors. This ensures that the model can account for color-based grouping in tidiness preferences.
B.2 Arrangement Strategy and Policy Setup
To generate a knolling dataset, a meticulous arrangement framework is developed, which integrates parameterized object properties, multiple preferences, and an iterative layout optimization process. This structured dataset generation ensures that the knolling model captures diverse object configurations, achieving precise arrangements. The framework enables training on a fixed number of objects (e.g., 10 objects), while flexibility in handling arbitrary numbers during deployment is achieved through masking techniques.
The knolling framework begins by setting arrangement policies that define the characteristics and layout behavior of each object. These policies include:
Object Dimensions: Length, width, and height ranges for each object are specified. For example, the length range is set to [0.036, 0.06] m, the width range to [0.016, 0.036] m, and the height range to [0.01, 0.02] m. This range allows for a controlled variation in object sizes within a standard workspace.
Classification Criteria: Objects are grouped based on attributes like type and color. Each classification criterion adds to the combinatorial complexity, as multiple preferences, creating distinct layouts can group objects.
Preferences and Policies: Policies for even distribution, forced alignment, and iteration constraints are set to control spatial relationships. These preferences guide the arrangement structure, affecting how objects are spaced, oriented, and aligned in the final layout.
This policy setup defines a complex framework where each additional preference increases layout complexity. Expanding beyond ten objects or adding more preferences exponentially intensifies the dataset generation’s computational requirements.
B.3 Arrangement Iteration
Object data, such as length, width, height, class, and color, are generated randomly, guided by the predefined limits set in the arrangement policies. With initial object attributes and groupings established, an iterative process is applied to optimize object layout. Iterations focus on minimizing the space occupied while meeting arrangement constraints, such as spacing, alignment, and rotational preferences. For each iteration, factors of the object count are calculated to determine potential grid configurations. For example, an object count of 10 allows configurations like a 2x5 grid, which are systematically tested. Objects are arranged iteratively in each grid configuration. During each iteration, the layout is scored based on spatial minimization and adherence to the arrangement preferences. Across iterations, the layout with the minimal space occupation and best policy adherence is selected as the final arrangement. Iteration count determines the thoroughness of this search, with higher values leading to more refined arrangements.
To improve computational efficiency, a threshold-based early-stopping mechanism is integrated. A threshold score for spatial occupation is predefined within the policy. If a layout configuration meets or surpasses this threshold, the iteration process halts early, reducing the total number of iterations. This early stopping criterion is crucial in managing computational resources, especially as the number of objects and arrangement preferences increases.
Appendix C Comparison of Knolling Bot and Related Methods
In this section, we compare our Knolling Bot approach with other notable methods in robotic object rearrangement, specifically StructFormer and LLM-GROP. A visual comparison is provided in Fig. 6. Our method learns generalizable tidiness patterns through a self-supervised learning framework, enabling it to rearrange an arbitrary number of objects into neat and visually organized layouts. This approach allows the robot to autonomously organize objects based on attributes such as color, size, or category, resulting in aesthetically pleasing and space-efficient arrangements. StructFormer employs a transformer-based neural network that interprets structured language commands and partial-view point clouds to rearrange objects into specified configurations. While effective in executing instructions like forming circles or lines, its reliance on explicit commands may limit its adaptability to more complex or unstructured tidiness preferences. Additionally, StructFormer’s focus on predefined structures may not fully capture the nuances of human-like organization, and its scalability to larger or more diverse object sets has not been extensively demonstrated. LLM-GROP leverages large language models (LLMs) combined with task and motion planning to achieve semantically valid object rearrangements in tabletop scenarios. It utilizes LLMs to extract commonsense knowledge about object configurations and integrates this with a task-motion planner to adapt to varying scene geometries. However, as demonstrated in its experiments, LLM-GROP is designed for tasks involving a limited number of objects, typically three to five, and focuses on specific tasks such as setting a dinner table. In contrast, Knolling Bot’s self-supervised learning framework enables it to handle a larger and more variable number of objects without explicit instructions, making it more versatile in organizing diverse and unstructured environments. Notably, Knolling Bot achieves a level of neatness and organization that closely aligns with human aesthetic preferences, arranging objects in a manner that is both visually pleasing and functionally efficient. This human-like understanding of tidiness distinguishes Knolling Bot from other methods
Figure 6: Comparison between our method and other methods. Our model can rearrange an arbitrary number of objects based on color, size, or categories and provide neat layouts.
Appendix D Gaussian Mixture Model for Multi-target Learning
$p(S_{i})=\sum_{j=1}^{J}\pi_{ij}\mathcal{N}(S_{i}|\mu_{ij},\sigma_{ij}^{2})$ (6)
Appendix E Position Encodering
Our model utilizes two position encoding methods to process the input data effectively. The first method, aligned with the original transformer paper’s approach, assigns unique tokens to each position in the list. This gives the data a sense of order, enabling the model to predict the target position of the initial objects and apply autoregression for the remaining objects. The second position encoding method maps the input data into a higher-dimensional space via sinusoidal functions. This method allows the model to handle higher frequency representations that potentially improve performance. The input, originally a 2-dimensional vector representing object length and width, is transformed into a 21-dimensional feature vector through this encoding method. By varying the wavelength at five different frequencies, the position encoding captures fine-grained patterns in the data, leading to more accurate predictions.
Appendix F Transformer Performance across Dataset Sizes
This experiment aimed to evaluate the performance of our knolling model across different scales of data. Four dataset sizes were selected for this study: 125k, 250k, 500k, and 1M. Each variant of the knolling model was trained independently until convergence. Following training, all models were evaluated using a consistent test dataset to maintain uniformity in the assessment.
The test errors obtained for each dataset size are shown in Table 5. The results highlight a consistent decline in error as the dataset size augments. This trend validates the widely accepted belief that transformers excel with the increase in data volume.
Appendix G Visual Perception Model Evaluation
G.1 YOLO Segmentation Model
Our visual perception module employs a customized YOLO v8 model, fine-tuned for object detection, segmentation, and attribute prediction. We trained and evaluated two versions of the model: one for the simulated environment (YOLO-sim), and another for real-world deployment, (YOLO-real).
G.1.1 Training Settings
Simulation Model (YOLO-sim):
  • Number of categories: 9
  • Number of objects per image: 4-11
  • Dataset sizes: 3200 train images, 800 test images
  • Pre-trained model: yolov8n-seg
Real-World Model (YOLO-real):
  • Number of categories: 9
  • Number of objects per image: 4-11
  • Dataset sizes: 80 train images, 20 test images
  • Pre-trained model: YOLO-sim
G.1.2 Evaluation Metrics and Results
We evaluated the models’ performance using the following metrics:
  • Average Accuracy: Measures the accuracy of category predictions.
  • Confusion Matrix: Provides a detailed breakdown of the model’s category prediction performance.
  • Average Hamming Distance: Assesses the quality of the predicted segmentation masks compared to the ground truth.
The evaluation results for both models are summarized in Tab.1.
Table 1: YOLO Segmentation Model Evaluation Results
Metric YOLO-sim YOLO-real
Average Hamming Distance 0.06743 0.05831
Average Accuracy 0.999 0.9857
Figure 7: Confusion matrix for the YOLO segmentation model. The matrix shows the model’s category prediction performance, with each cell representing the number of instances predicted as a particular category (column) compared to the true category (row). The diagonal cells indicate correct predictions, while off-diagonal cells represent misclassifications.
The results demonstrate that both models achieve high accuracy in category prediction and generate precise segmentation masks, with the real-world model slightly outperforming the simulation model in terms of Average Hamming Distance. The confusion matrices for both models are provided in Fig.7.
These evaluation results validate the effectiveness of our visual perception module in accurately detecting, segmenting, and categorizing objects in both simulated and real-world environments, enabling robust input for the subsequent knolling pipeline.
Appendix H Comparison to Baselines
To further validate the effectiveness of our transformer-based knolling model, we conducted additional experiments, exploring the model’s performance relative to widely used architectures, including MLP, LSTM, and CNN2d (Table2) . These comparisons go beyond the strict parameter-matching paradigm outlined in the main paper, allowing architectural latitude in the baseline models to showcase their maximum potential performance under flexible conditions.
The motivations for this experiment are twofold. First, it addresses the practical concern that simpler models, such as MLPs or LSTMs, may perform comparably with smaller datasets. By demonstrating the superior performance of the transformer-based model even when baselines are given additional parameters, we underscore its robustness and suitability for complex layout tasks like knolling. Secondly, this comparison contributes to understanding the limitations of conventional architectures in capturing spatial dependencies critical to knolling applications.
Our transformer-based model significantly outperforms these baselines, where we report the error and parameter count for each model and configuration. This comparison illustrates that even with increased parameters, the baseline architectures exhibit higher error rates, reinforcing the advantages of a transformer-based approach for Knolling task.
Table 2: Comparison to more baselines
MLP Error Parameters LSTM Error Parameters CNN2d Error Parameters
1 3.73E-01 62079 1 3.15E-02 183732 1 3.08E-02 92930
2 3.46E-01 107205 2 3.06E-02 101740 2 3.01E-01 92930
3 3.20E-01 10885 3 3.14E-02 73152 3 3.99E-01 40081
4 3.25E-01 125540 4 2.70E-02 191480 4 3.95E-01 42520
5 1.49E-01 101979 5 2.91E-02 103236 5 4.00E-01 36941
6 2.98E-01 136815 6 2.57E-02 60622 6 3.62E-01 30344
7 1.40E-01 83100 7 3.15E-02 62372 7 3.42E-01 79924
8 2.32E-01 8425 8 2.91E-02 81524 8 4.67E-01 27620
9 2.07E-01 70335 9 3.09E-02 90990 9 4.01E-01 34670
10 3.41E-01 170025 10 2.32E-02 183732 10 3.93E-01 33196
MEAN 2.73E-01 MEAN 2.90E-02 MEAN 3.49E-01
STD 8.02E-02 STD 2.69E-03 STD 1.14E-01
MIN 1.40E-01 MIN 2.32E-02 MIN 3.08E-02
Appendix I Quantitative Evaluation and Ablation Study
@@
Table 3: Comparison between Our Method (OM) and Baselines (LSTM, MLP)
@@
Table 4: Ablation study
@@
Two baseline models were established to enable a performance comparison to Our Method (OM). The first baseline model utilized a Multilayer Perceptron (MLP) architecture, and the second baseline was designed based on the Long Short-Term Memory (LSTM) model. We use LSTM as a baseline due to its effectiveness in sequence prediction tasks. Arranging objects can be framed as predicting a sequence of positions, where each object’s placement is influenced by prior placements. LSTMs have been extensively used for various sequence prediction tasks and have shown good performance. Thus, LSTMs serve as a more challenging baseline for benchmarking the performance of our transformer-based model. Moreover, comparing Transformer model with LSTMs helps to highlight the benefits of using self-attention and auto-regression in the transformer for handling variable input and output sizes and multi-label problems. Just as we did with Transformer model, we measured the L1 distance between the actual and predicted positions for each baseline model, with the results detailed in Table 3. As a commitment to a fair evaluation, we have ensured that each model utilizes a similar amount of parameters: Transformer model incorporates 87,458 parameters, the LSTM baseline uses 86,858 parameters, and the MLP baseline operates with 87,788 parameters. Our knolling model consistently outperforms the MLP and LSTM baselines in terms of the L1 error. This superiority of Transformer model is evident in all parameters: mean L1 error, standard deviation, and minimum and maximum error. These results validate the effectiveness of our transformer-based model, which employs self-attention and auto-regression in handling varying input sizes and multi-label problems in knolling tasks.